OpenStudy (anonymous):

I just did an integral and i'm trying to simplify my final answer: xln(2x+1)-x+(1/2)(ln(2x+1))+C

OpenStudy (mr.math):

Is it \(x\ln(2x+1)-x+\frac{1}{2}\ln(2x+1)+C?\)

OpenStudy (anonymous):

yes

OpenStudy (mr.math):

I would recommend that you take \(\ln(2x+1)\) as a common factor.

OpenStudy (mr.math):

Could you tell me what the original integral was?

OpenStudy (anonymous):

the problem is that the answer is \[(1/2)(2x+1)\ln(2x+1)-x+C\] What happened to the x in xln(2x+1)

OpenStudy (anonymous):

The original integral was [int_ln(2x+1)dx]

OpenStudy (anonymous):

i set u=ln(2x+1) and dv=dx

OpenStudy (mr.math):

Your answer is right then. Here how they did it: \[x\ln(2x+1)+\frac{1}{2}\ln(2x+1)-x+C=(x+\frac{1}{2})\ln(2x+1)-x+C\]

OpenStudy (mr.math):

But if you take \(\frac{1}{2}\) as a common factor from the first parentheses, you get \[\frac{1}{2}(2x+1)\ln (2x+1)-x+C\]

OpenStudy (anonymous):

I see. So you multiply by 2 within the parenthesis and divide by 2 to make up for the difference

OpenStudy (mr.math):

Exactly!

OpenStudy (anonymous):

Ur awesome. Thanks again my friend

OpenStudy (mr.math):

You learn fast by the way :)

OpenStudy (anonymous):

I disagree...but I'll take the compliment

OpenStudy (mr.math):

You have to :P